1)15x-2-14x+8=13
2)-5(3x+4)+10=25
3)5(y-6)+(y+2)=44

Answer by CubeyThePenguin(3113)

1) 15x - 2 - 14x + 8 = 13
x = 7

2) -5(3x + 4) + 10 = 25
-5(3x + 4) = 15
3x + 4 = -3
3x = -7
x = -7/3

3) 5(y-6) + (y-2) = 44
5y-6 + y-2 = 44
6y = 52
y = 26/3